OpenStudy (anonymous):

Simpliest version of 4x/4(x+y)...is the answer x/(x+y) or not??? Will give medalssss

OpenStudy (anonymous):

Please help

jimthompson5910 (jim_thompson5910):

yes, \[\Large \frac{4x}{4(x+y)} = \frac{x}{x+y}\] where x and y are nonzero and x does not equal -y. This is all to avoid dividing by zero

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

sure thing